Circular No. 8420

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2004fa AND 2004fb Further to IAUC 8407, Y.-t. Chen et al. (plus W. Hsiao) report the

discovery of two apparent supernovae on unfiltered CCD images taken with the 1-m Lulin Observatory telescope.

SN  2004 UT  α_{2000}  δ_{2000}  Mag.  Offset
2004fa Sept. 5.672 20^{h}22^{m}35.^{s}00 −29^{o}49′20′′.7 16.8 0′′.2 W, 4′′.3 N
2004fb Oct. 7.698 20 15 25.63 −37 30 34.6 18.3 1′′.7 E, 7′′.1 N

Additional approximate magnitudes: SN 2004fa in MCG –05-48-5 = PGC 64555, Aug. 13.591 UT, [19.3; 30.65, 19.1; Sept. 1.536, 18.1; 2.608, 17.8. SN 2004fb in ESO 340-G7, Sept. 1.699, [19.0; Oct. 9.675, 18.5; 10.694, 18.5.

SUPERNOVAE 2004dx, 2004eh, 2004ep, 2004eq, 2004ev, 2004ex–2004ez

A. V. Filippenko, R. J. Foley, and R. Chornock, University of California, Berkeley; and T. Matheson, National Optical Astronomy Observatory, report that inspection of CCD spectra (range 320–940 nm), obtained on Oct. 18 UT with the Keck I 10-m telescope (+ LRIS), shows that SN 2004dx (IAUC 8392) is a supernova, probably of type Ic (with a remote chance of being of type Ib), 2–3 months past maximum brightness. SN 2004eh (IAUC 8401) is of type Ib, roughly a month past maximum; the He I lines have well-developed P-Cyg profiles. SN 2004ep (IAUC 8406) is of type II in the nebular phase, many months past maximum. SN 2004eq (IAUC 8407, 8411) is of type Ia, ∼ 2 months past maximum. SN 2004ev (IAUC 8417) is also of type Ia, perhaps 3 weeks past maximum. SN 2004ex (IAUC 8418) is of type II, within a few weeks past maximum; the continuum is relatively blue, Hα has a classic P-Cyg profile, and the other Balmer and Fe II lines are strong. The ejecta velocity measured from the Hα absorption minimum is unusually large, ∼ 14000 km/s. SN 2004ey (IAUC 8419) is a young, normal type-Ia supernova close to maximum brightness. SN 2004ez (IAUC 8419) is of type II, perhaps 1–2 months past maximum, with prominent P-Cyg profiles of hydrogen, Fe II, and Ca II; the ejecta velocity measured from the Hα absorption minimum is ∼ 4600 km/s.

SUPERNOVA 2004dj IN NGC 2403 CCD R-band magnitudes by K. Hornoch, Lelekovice, Czech Republic:

Sept. 4.976 UT, 11.71; 7.001, 11.72; 14.033, 11.86; 17.932, 11.88; 30.941, 12.23; Oct. 11.892, 12.77.

2004 October 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green